OpenStudy (jiteshmeghwal9):

If sum of the roots is \(p\) & the sum of their square is \(q^2\) the equation is :-

hartnn (hartnn):

hint : \((a+b)^2=a^2+b^2+2ab\) you can find product of roots from here.

hartnn (hartnn):

a+b=p a^2+b^2=q^2

OpenStudy (jiteshmeghwal9):

sorry but i couldn't understand it very well :(

OpenStudy (jiteshmeghwal9):

only gt u a little :(

OpenStudy (hba):

Consider 'a' and 'b' are the roots of the equation.

OpenStudy (hba):

So their sum is p a+b=p

hartnn (hartnn):

you got the equation if 1st comment, right ? substituting the values. \(ab=(p^2-q^2)/2=product \: of \: roots\)

hartnn (hartnn):

now u have sum and product, can u find the equation ??

hartnn (hartnn):

\(x^2-sum*x+product=0\)

OpenStudy (hba):

\[x^2-(S.O.R)x+(P.O.R)=0\]

OpenStudy (jiteshmeghwal9):

ohh k i gt it

OpenStudy (jiteshmeghwal9):

Thanx both of u a lot :)

hartnn (hartnn):

welcome ^_^

OpenStudy (hba):

Welcome :)